OpenStudy (anonymous):

Create your own binomial expression with a radical in the second term. Part 1: Identify its conjugate and explain, in complete sentences, why it is the conjugate. (1 point) Part 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why? (1 point

OpenStudy (anonymous):

x-sqrt2 conjugate = x + sqrt2 (x-sqrt2)(x+sqrt2) = x^2 -xsqrt2 +xsqrt2 - 2 = x^2 - 2 the radicals are eliminated during the multiplaction